htm Chapter 4 Assignment Help Save & Ex 7 Assume that licence plates have leters followed by numbers a. Some Saskatchewan licence plates have three letters and three numbers. How l many licence plates of this type can there be? b. Some Ontario licence piates have four lethers and three numbers How many licence plates of this type can there be? References
Answer #1

There are 10 possible digits for the numbers (0,1,2,...,9), and 26 possible letters (A,B,C,.....,Z).

Since repetitions are allowed, we have that for each letter used, 26 still remain for the next choice, and for each digit used, 10 still remain for the next choice.

a) Since the Saskatchewan license have three letters and three numbers.

Hence, by the multiplication principle, the total different number of letter permutations is 26×26×26=17576 and the total number of digit permutations is 10×10×10=1000.

Thus the number of Saskatchewan license plate is 17576 × 1000= 17576000

b) Since ontario license have four letters and three numbers

Hence, by the multiplication principle, the total different number of letter permutations is 26×26×26×26=456976 and the total number of digit permutations is 10×10×10= 1000.

Thus the number of ontario license plate is 456976 × 1000= 456976000